# Semiclassical measure of the propagation between two topological insulators

### v2

@@ Appendix C Spectral Analysis of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$
$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)=\sum_{n\in\mathbb{Z}}\lambda_{n}(s,\sigma)\Pi_{n}(s,\sigma).$
Proof of Theorem C.1.
First, we will rescale the function in order to have the form $f\left(s,r(s)y\right)$. We can do it because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non negative (Assumption 1.2). Then we dilate with $r(s)^{-1}$ and we denote by $\zeta$ the ratio $\sigma/r(s)$. For all $f\in\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
$\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Bigg{[}f\Big{(}r(s)y\Big{)}\Bigg{]}$ $=\left[\begin{pmatrix}\sigma&r(s)(y+\mathrm{i}\mathrm{D}_{y})\\ r(s)(y-\mathrm{i}\mathrm{D}_{y})&-\sigma\end{pmatrix}f\right]\Big{(}r(s)y\Big{)}$
$=r(s)\left[\begin{pmatrix}\zeta&y+\mathrm{i}\mathrm{D}_{y}\\ y-\mathrm{i}\mathrm{D}_{y}&-\zeta\end{pmatrix}f\right]\Big{(}r(s)y\Big{)}.$
Let us define
$\mathscr{T}\coloneqq\begin{pmatrix}\zeta&\mathfrak{a}\\ \mathfrak{a}^{\dagger}&-\zeta\end{pmatrix},$
where $\mathfrak{a}\coloneqq y+\mathrm{i}\mathrm{D}_{y}$ denotes the usual annihilation operator on the variable $y$. Then the link between the spectrum is
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=r(s)\mathrm{Sp}(\mathscr{T})$
So we can study $\mathscr{T}$ instead of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$. Let $\lambda_{\zeta}\in\mathrm{Sp}(\mathscr{T}),\begin{pmatrix}f\\ g\end{pmatrix}\in E_{\lambda_{\zeta}}(\mathscr{T})$ if and only if
$\left\{\begin{matrix}2\zeta f&+&\mathfrak{a}g&=&\left(\lambda_{\zeta}+\zeta\right)f\\ &&\mathfrak{a}^{\dagger}f&=&\left(\lambda_{\zeta}+\zeta\right)g\end{matrix}\right.$
If $\lambda_{\zeta}=-\zeta$, the injectivity of $\mathfrak{a}^{\dagger}$ on $\mathrm{L}^{2}\left(\mathbb{R}\right)$ implies $f=0$ whence $\mathfrak{a}g=0$ so $g=A\mathfrak{h}_{0}$ for some $A\in\mathbb{C}$. So $-\zeta\in\mathrm{Sp}(\mathscr{T})$ and is a simple eigenvalue with ${\begin{pmatrix}0\\ \mathfrak{h}_{0}\end{pmatrix}}$ a normalized eigenfunction. If $\lambda_{\zeta}\neq-\zeta$, then
$\left\{\begin{matrix}\mathfrak{a}\mathfrak{a}^{\dagger}f&=&\left(\lambda_{\zeta}^{2}-\zeta^{2}\right)f\\ g&=&\frac{1}{\lambda_{\zeta}+\zeta}\mathfrak{a}^{\dagger}f\end{matrix}\right.$
and eigenvalues of $\mathfrak{a}\mathfrak{a}^{\dagger}$ are known so we deduce that there exists $n\in\mathbb{N}_{>0},\lambda_{\zeta}^{2}-\zeta^{2}=2n$ so
$\lambda_{\zeta}\in\left\{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}\ |\ n\in\mathbb{Z}_{\neq 0}\right\}=\mathrm{Sp}\left(\mathfrak{a}\mathfrak{a}^{\dagger}\right),$
@@
$\forall(\zeta,y)\in\mathbb{R}^{2},\ \begin{pmatrix}f_{n}\\ g_{n}\end{pmatrix}(\zeta,y)\coloneqq B\begin{pmatrix}\mathfrak{h}_{|n|-1}(y)\\ \frac{\sqrt{2|n|}}{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}+\zeta}\mathfrak{h}_{|n|}(y)\end{pmatrix}.$
@@
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\}.$
The associated projectors $\left(\Pi_{n}\right)_{n\in\mathbb{Z}}$ are smooth because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non negative (Assumption 1.2); for all $n\in\mathbb{Z}_{\neq 0}$, $\lambda_{n}$ and $\lambda_{n}+\sigma$ is smooth and never vanishes. ∎

### v3

@@ Appendix B Spectral Analysis of the principal normal form operator
$\mathrm{T}_{\mathrm{E}}(s,\sigma)=\sum_{n\in\mathbb{Z}}\lambda_{n}(s,\sigma)\Pi_{n}(s,\sigma).$
Proof.
According to Remark B.1, we study $\mathscr{T}$. Let $\lambda_{\zeta}\in\mathrm{Sp}(\mathscr{T}),\begin{pmatrix}f\\ g\end{pmatrix}\in E_{\lambda_{\zeta}}(\mathscr{T})$ if and only if
$\left\{\begin{matrix}&&\mathfrak{a}^{\dagger}g&=&\left(\lambda_{\zeta}-\zeta\right)f\\ \mathfrak{a}f&-&2\zeta g&=&\left(\lambda_{\zeta}-\zeta\right)g\end{matrix}\right$.
If $\lambda_{\zeta}=\zeta$, then the injectivity of $\mathfrak{a}^{\dagger}$ on $\mathrm{L}^{2}\left(\mathbb{R}\right)$ implies $g=0$ whence $\mathfrak{a}f=0$ so $f=A\mathfrak{h}_{0}$ for some $A\in\mathbb{C}$. So $\zeta\in\mathrm{Sp}(\mathscr{T})$ and is a simple eigenvalue with ${\begin{pmatrix}\mathfrak{h}_{0}\\ 0\end{pmatrix}}$ a normalized eigenfunction. If $\lambda_{\zeta}\neq\zeta$, then
$\left\{\begin{matrix}f&=&\frac{1}{\lambda_{\zeta}-\zeta}\mathfrak{a}^{\dagger}g\\ \mathfrak{a}\mathfrak{a}^{\dagger}g&=&\left(\lambda_{\zeta}^{2}-\zeta^{2}\right)g\end{matrix}\right$.
so, according to Proposition B.2, we deduce that there exists $n\in\mathbb{N}_{>0},\lambda_{\zeta}^{2}-\zeta^{2}=2n$ so
$\lambda_{\zeta}\in\left\{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}\ |\ n\in\mathbb{Z}_{\neq 0}\right\},$
@@
$\forall(\zeta,y)\in\mathbb{R}^{2},\ \begin{pmatrix}f_{n}\\ g_{n}\end{pmatrix}(\zeta,y)\coloneqq B\begin{pmatrix}\frac{\sqrt{2|n|}}{\mathrm{sgn}(n)\sqrt{\zeta^{2}+2|n|}-\zeta}\mathfrak{h}_{|n|}(y)\\ \mathfrak{h}_{|n|-1}(y)\end{pmatrix}$.
@@
$\mathrm{Sp}\Big(\mathrm{T}_{\mathrm{E}}(s,\sigma)\Big)=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\}.$
The associated projectors $\left(\Pi_{n}\right)_{n\in\mathbb{Z}}$ are smooth because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non-negative (Assumption 1.1); for all $n\in\mathbb{Z}_{\neq 0}$, $\lambda_{n}$ and $\lambda_{n}+\sigma$ is smooth and never vanishes. ∎
Acknowledgments
The author is very grateful to his advisors Clotilde Fermanian-Kammerer and Nicolas Raymond for many insightful discussions and valuable suggestions. The author acknowledges the support of the Région Pays de la Loire via the Connect Talent Project HiFrAn 2022 07750, and from the France 2030 program, Centre Henri Lebesgue ANR-11-LABX-0020-01.